A harvard english professor attempts to estimate the number of novels that the students read during their time in college. professor expects to have margin error of five books at 80% confidence & expects to have a deviation of 15 books to read. how many students should he survey?
Answer #1

Sample size n = (Z\alpha/2 * \sigma / E)2

Where,

Z\alpha/2 is critical value

\sigma is standard deviation

E is margin of error

n = (1.2816 * 15 / 5)2

= 14.78

= 15 (Rounded up to nearest integer)

Number of students should be survey = 15